Help to solve and figure out Rose's garden is in the shape of a trapezoid. If the height of the trapezoid is 16m, one base is 20m, and the area is 224m2, i need t find the length of the other base . Word problems always gave me a hard time I really could not under stand them
Answer by checkley71(8403)

(20+X)/2*16=224
(20+X)8=224
160+8X=224
8X=224-160
8X=64
X=64/8
X=8 THE OTHER SIDE.
PROOF
28/2*16=224
28*8=224
224=224